OpenStudy (anonymous):

QUESTION FOR DUMMIES & intelligent ppl can also try it Question (physics + math) A ship is moving at a speed of 56 km/h. One second later, It's moving at 58 km/h. Whats it's accelaration..

OpenStudy (anonymous):

@baddinlol

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

hmm..... Solve it then

OpenStudy (anonymous):

Are you familiar with newtons constant acceleration equations?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

v=u+at

OpenStudy (anonymous):

u = 56/3.6 = 15.55m/s

OpenStudy (lgbasallote):

\[v_2 = v_1 + at\] \[58 = 56 + a\] \[58 - 56 = a\]

OpenStudy (anonymous):

okay!! then

OpenStudy (anonymous):

@lgbasallote Plzz try to give a good solution by briefing each steps

OpenStudy (lgbasallote):

was that not steo-by-step enough

OpenStudy (anonymous):

 RIGHT